OpenStudy (anonymous):

Fluctuation in the prices of precious metals such as gold have been empirically shown to be well approximated by a normal distribution when observed over short interval of time. In May 1995, the daily price of gold (1 troy ounce) was believed to have a mean of $383 and a standard deviation of $12. A broker, working under these assumptions, wanted to find the probability that the price of gold the next day would be between $394 and $399 per troy ounce. In this eventuality, the broker had an order from a client to sell the gold in the client's portfolio. What is the probability that the client's

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

Find the z-scores of the two numbers. Use a z-table to find the probabilities of less than each z-score. Subtract the probability of the 394 score from the 399 one. There's your answer.

OpenStudy (anonymous):

I know the answers in 9% just need to see how one gets to that answer.

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

I need to see the numbers used in the z tables to show me how you get to the answer. I'm new to all this.

OpenStudy (anonymous):

http://allianthawk.org/images/z_table_positive_bbb.png Use the z-formula (z=x-mean/stddev) to find the two z-values you need.

OpenStudy (anonymous):

For example, z of 394 = 394-383/12 = 0.917 = approx 0.92. The z-value for 0.92 is .8212.

OpenStudy (anonymous):

Do you know how to read a z-table?

OpenStudy (anonymous):

I see on the chart where the .92= .8212

OpenStudy (anonymous):

Use that formula to find the z-score of 399. (z= (x-mean)/stddev) What value do you get from that z-score?

OpenStudy (anonymous):

z(399) = 399-383 / 12 = 1.3333

OpenStudy (anonymous):

Ok, now what percentage do you see that corresponds to 1.33 in the table?

OpenStudy (anonymous):

.9066

OpenStudy (anonymous):

So the probability of the price being between 394 and 399 is .9066-.8212 = .0854

OpenStudy (anonymous):

I see it now... THANK YOU!!!